without actually solving the equation, list all possible numbers that would be rejected if they appeared as potential solutions. 1/x+5 - 1/x-5= 0
Found 2 solutions by amoresroy, CharlesG2:
Answer by amoresroy(333)

without actually solving the equation, list all possible numbers that would be rejected if they appeared as potential solutions. 1/x+5 - 1/x-5= 0
The possible numbers are -5 and 5. These values of x will make the denominator equal to zero.

Answer by CharlesG2(828)

without actually solving the equation, list all possible numbers that would be rejected if they appeared as potential solutions. 1/x+5 - 1/x-5= 0

if this is 1/x + 5 - 1/x - 5 = 0, the 1/x would cancel out so would the 5,
leaving 0 = 0, so would be nothing to solve for

if this is 1/(x + 5) - 1/(x - 5) = 0
5 and -5 would be rejected since can not divide by zero
(x - 5)/[(x + 5)(x - 5)] - (x + 5)/[(x - 5)(x + 5)] = 0
x - 5 - (x + 5) = x - 5 - x - 5 = -10
-10/(x^2 - 25) = 0
there is no way -10 divided by anything is going to be zero,
there is no solutions